OpenStudy (anonymous):

108 is 36% of what number? Write and solve the proportion to solve the problem.

OpenStudy (luigi0210):

Try this: \[\LARGE \frac{36}{100}=\frac{108}{x} \]

OpenStudy (anonymous):

108/36=3 3*100=300 300

OpenStudy (anonymous):

If you divide the number by the percentage it gives you the value of 1%, if you then multiply that by 100 it is the value of 100%

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Yes thanks!

OpenStudy (texaschic101):

is/of = %/100 ==> percent formula 108 IS 36% OF x 108/x = 36/100 cross multiply 36x = (100)(108) 36x = 10800 x = 10800/36 x = 300